# Mamba-FSCIL: Dynamic Adaptation with Selective State Space Model for Few-Shot Class-Incremental Learning

### v1

@@ 5.1 Datasets and Implementation Details
Self-promoted [82] 64.10 65.86 61.36 57.45 53.69 50.75 48.58 45.66 43.25 54.52 +14.26
Data-free Replay [29] 74.40 70.20 66.54 62.51 59.71 56.58 54.52 52.39 50.14 60.78 +7.37
MetaFSCIL [9] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97 60.79 +7.54
C-FSCIL [22] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47 61.64 +7.04
LIMIT [81] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23 61.84 +6.28
ALICE [42] 79.00 70.50 67.10 63.40 61.20 59.20 58.10 56.30 54.10 63.21 +3.41
CABD [78] 79.45 75.38 71.84 67.95 64.96 61.95 60.16 57.67 55.88 66.14 +1.63
NC-FSCIL [70] 82.52 76.82 73.34 69.68 66.19 62.85 60.96 59.02 56.11 67.50 +1.40
@@
TOPIC [57] 64.10 55.88 47.07 45.16 40.11 36.38 33.96 31.55 29.37 42.62 +28.14
CEC [76] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14 59.53 +8.37
DSN [68] 73.00 68.83 64.82 62.24 59.16 56.96 54.04 51.57 49.35 60.00 +8.16
Mamba-FSCIL (ours) 82.89 77.85 73.69 69.67 66.89 63.66 61.48 59.74 57.51 68.12
Table 3: Performance of FSCIL in each session on CUB-200 and comparison with other studies.“Average Acc.” is the average accuracy of all sessions. “Final Improv.” calculates the improvement of our method in the last session over previous methods.
Methods Accuracy in each session (%) $\uparrow$ Average Final
@@
Fixed classifier [43] 68.47 51.00 45.42 40.76 35.90 33.18 27.23 24.24 21.18 17.34 16.20 34.63 +45.45
Self-promoted [82] 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33 49.32 +24.32
SFbFSCIL [8] 68.78 59.37 59.32 54.96 52.58 49.81 48.09 46.32 44.33 43.43 43.23 51.84 +18.42
DeepInv [73] 75.90 70.21 65.36 60.14 58.79 55.88 53.21 51.27 49.38 47.11 45.67 57.54 +15.98
D-DeepEMD [75] 75.35 70.69 66.68 62.34 59.76 56.54 54.61 52.52 50.73 49.20 47.60 58.73 +14.05
D-NegCosine [28] 74.96 70.57 66.62 61.32 60.09 56.06 55.03 52.78 51.50 50.08 48.47 58.86 +13.18
Data-free replay [29] 75.90 72.14 68.64 63.76 62.58 59.11 57.82 55.89 54.92 53.58 52.39 61.52 +9.26
MetaFSCIL [9] 75.9 72.41 68.78 64.78 62.96 59.99 58.3 56.85 54.78 53.82 52.64 61.93 +9.01
MgSvF [77] 72.29 70.53 67.00 64.92 62.67 61.89 59.63 59.15 57.73 55.92 54.33 62.37 +7.32
FACT [80] 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94 64.42 +4.71
IDLVQ [5] 77.37 74.72 70.28 67.13 65.34 63.52 62.10 61.54 59.04 58.68 57.81 65.23 +3.84
LIMIT [81] 76.32 74.18 72.68 69.19 68.79 65.64 63.57 62.69 61.47 60.44 58.45 66.67 +3.20
NC-FSCIL [70] 80.45 75.98 72.30 70.28 68.17 65.16 64.43 63.25 60.66 60.01 59.44 67.28 +2.21
ALICE [42] 77.40 72.70 70.60 67.20 65.90 63.40 62.90 61.90 60.50 60.60 60.10 65.75 +1.55
@@
TOPIC [57] 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28 43.92 +35.37
CEC [76] 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28 61.33 +9.37
DSN [68] 76.06 72.18 69.57 66.68 64.42 62.12 60.16 58.94 56.99 55.10 54.21 63.31 +7.44
Mamba-FSCIL (ours) 80.90 76.26 72.97 70.14 67.83 65.74 65.43 64.12 62.31 62.12 61.65 68.22
5.2 Comparison with State-of-the-art Methods
The comparison with the state-of-the-art methods on minImageNet, CIFAR-100, and CUB-200 is shown in Tab. 1, Tab. 2, and Tab. 3. The results demonstrate that Mamba-FSCIL consistently outperforms existing static and dynamic adaptation methods by over 1% in terms of the last-session accuracy, where all classes are taken into account. Specifically, on miniImageNet, Mamba-FSCIL achieves the highest average accuracy of 69.81% and increases the last-session accuracy to 59.36%, which surpasses the challenging baseline NC-FSCIL [70] by 1.99% and 1.05%, respectively. In CIFAR-100, Mamba-FSCIL maintains the best performance in most sessions, with the last session accuracy outperforming NC-FSCIL by 1.4%. Notably, Mamba-FSCIL significantly surpasses dynamic methods like DSN [68] with an average accuracy improvement of 8.12% over DSN. In the more challenging fine-grained CUB-200 dataset, Mamba-FSCIL leads with an average accuracy of 68. 22%, outperforming all static and dynamic methods. Compared to DSN, Mamba-FSCIL shows the remarkable ability to learn complex classification scenarios incrementally, improving the accuracy in the last session by 7.44%, also indicating a lower performance drop over sessions.
@@ 5.3 Ablation Studies
Table 4: Ablation studies to investigate the effects of our dual selective projector and the class-sensitive selective scan mechanism on FSCIL performance across different datasets. “Base” refers to the accuracy of the first session; “Final” refers to the last session; “PD” denotes the performance drop, i.e., the accuracy difference between the first and the last session.
Methods Adaptation Type miniImageNet CIFAR-100 CUB-200
@@
In our ablation studies across the three datasets, we assess the impact of our dual selective projector and class-sensitive selective scan mechanism. Utilizing a consistent setup that includes the same backbone, identity branch, memory module, classifier, and classification loss, we explore several configurations: (1) Baseline employs NC-FSCIL with a static MLP-based projection layer, serving as a competitive static adaptation method. (2) + Single Selective SSM Projector integrates a single dynamic SSM projector with one projection branch, updated across all sessions. (3) + Dual Selective SSM Projector uses our dual selective SSM projector with two branches, as introduced in Sec. 4.2. (4) + Dual Selective SSM Projector + Class-sensitive Selective Scan further integrates our class-sensitive selective scan.
@@
The introduction of a single selective SSM projector improves base session accuracy across the three datasets (e.g., from 84.02% to 84.93% on miniImageNet and from 82.52% to 82.80% on CIFAR-100), demonstrating the Mamba-based projector helps to improve the performance in the base session, providing a more powerful feature extractor for the subsequent sessions. However, despite these gains, the final performances on miniImageNet and CIFAR-100 both fall below the baseline, implying that a naive replacement of the MLP projector with the selective SSM projector cannot effectively mitigate catastrophic forgetting, increasing the performance drop (PD) as a result. In contrast, when our proposed dual selective SSM projector is adopted, the final performance is also effectively improved over the baseline, reducing the performance drop from 27.04% to 26.01% on miniImageNet and from 26.73% to 26.09% on CIFAR-100. This shows that our dual design is capable of adaptively learning feature shifts of novel classes into their corresponding classifier prototypes with fewer disruptions to the learned knowledge of the base classes. Further improvements are achieved with the integration of our class-sensitive scan mechanism into the dual projector, increasing the last-session accuracy from 56.71% to 57.51% on CIFAR-100 and from 61.06% to 61.65% on CUB-200. These findings validate the critical role of each component in Mamba-FSCIL on multiple datasets.
We also conduct a sensitivity analysis on the hyper-parameters $\alpha$ and $\beta$ of Eq. (22) in Tab. 5 to evaluate their impact on our model’s average accuracy on miniImageNet. Each experiment varies one parameter with the other fixed. The results show that when $\alpha$ and $\beta$ are set within a proper range, the performance only varies slightly, demonstrating the model’s robustness to the hyper-parameters.
Table 5: Effects of hyper-parameters $\alpha$ and $\beta$.

### v2

@@ 5.2 Comparison with the State-of-the-art Methods
Self-promoted [86] CVPR 2021 64.10 65.86 61.36 57.45 53.69 50.75 48.58 45.66 43.25 54.52 +14.26
Data-free Replay [6] ECCV 2022 74.40 70.20 66.54 62.51 59.71 56.58 54.52 52.39 50.14 60.78 +7.37
MetaFSCIL [17] CVPR 2022 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97 60.79 +7.54
C-FSCIL [16] CVPR 2022 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47 61.64 +7.04
LIMIT [18] TPAMI 2022 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23 61.84 +6.28
FACT [3] CVPR 2022 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10 62.24 +5.41
TEEN [87] NeurIPS 2023 74.92 72.65 68.74 65.01 62.01 59.29 57.90 54.76 52.64 63.10 +4.87
ALICE [8] ECCV 2022 79.00 70.50 67.10 63.40 61.20 59.20 58.10 56.30 54.10 63.21 +3.41
CABD [88] CVPR 2023 79.45 75.38 71.84 67.95 64.96 61.95 60.16 57.67 55.88 66.14 +1.63
NC-FSCIL [19] ICLR 2023 82.52 76.82 73.34 69.68 66.19 62.85 60.96 59.02 56.11 67.50 +1.40
@@
TOPIC [1] CVPR 2020 64.10 55.88 47.07 45.16 40.11 36.38 33.96 31.55 29.37 42.62 +28.14
LEC-Net [21] arXiv 2022 64.10 53.23 44.19 41.87 38.54 39.54 37.34 34.73 34.73 43.14 +22.78
CEC [2] CVPR 2021 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14 59.53 +8.37
DSN [23] TPAMI 2022 73.00 68.83 64.82 62.24 59.16 56.96 54.04 51.57 49.35 60.00 +8.16
FeSSSS [89] CVPR 2022 75.35 70.81 66.70 62.73 59.62 56.45 54.33 52.10 50.23 60.92 +7.28
Mamba-FSCIL (ours) 82.80 77.85 73.69 69.67 66.89 63.66 61.48 59.74 57.51 68.14
TABLE III: FSCIL performance across sessions on CUB-200 compared with other methods. ”Average Acc.” is the average accuracy across all sessions. ”Final Improv.” shows the last session’s accuracy improvement of our method over previous methods.
Methods Venue Accuracy in each session (%) $\uparrow$ Average Final
@@
Self-promoted [86] CVPR 2021 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33 49.32 +24.32
SFbFSCIL [80] ICCV 2021 68.78 59.37 59.32 54.96 52.58 49.81 48.09 46.32 44.33 43.43 43.23 51.84 +18.42
Data-free replay [6] ECCV 2022 75.90 72.14 68.64 63.76 62.58 59.11 57.82 55.89 54.92 53.58 52.39 61.52 +9.26
MetaFSCIL [17] CVPR 2022 75.9 72.41 68.78 64.78 62.96 59.99 58.3 56.85 54.78 53.82 52.64 61.93 +9.01
MgSvF [116] TPAMI 2021 72.29 70.53 67.00 64.92 62.67 61.89 59.63 59.15 57.73 55.92 54.33 62.37 +7.32
FACT [3] CVPR 2022 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94 64.42 +4.71
IDLVQ [11] ICLR 2021 77.37 74.72 70.28 67.13 65.34 63.52 62.10 61.54 59.04 58.68 57.81 65.23 +3.84
LIMIT [18] TPAMI 2022 76.32 74.18 72.68 69.19 68.79 65.64 63.57 62.69 61.47 60.44 58.45 66.67 +3.20
TEEN [87] NeurIPS 2023 77.26 76.13 72.81 68.16 67.77 64.40 63.25 62.29 61.19 60.32 59.31 66.63 +2.34
NC-FSCIL [19] ICLR 2023 80.45 75.98 72.30 70.28 68.17 65.16 64.43 63.25 60.66 60.01 59.44 67.28 +2.21
ALICE [8] ECCV 2022 77.40 72.70 70.60 67.20 65.90 63.40 62.90 61.90 60.50 60.60 60.10 65.75 +1.55
@@
TOPIC [1] CVPR 2020 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28 43.92 +35.37
LEC-Net [21] arXiv 2022 70.86 58.15 54.83 49.34 45.85 40.55 39.70 34.59 36.58 33.56 31.96 45.08
CEC [2] CVPR 2021 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28 61.33 +9.37
FeSSSS [89] CVPR 2022 79.60 73.46 70.32 66.38 63.97 59.63 58.19 57.56 55.01 54.31 52.98 62.85 +8.67
DSN [23] TPAMI 2022 76.06 72.18 69.57 66.68 64.42 62.12 60.16 58.94 56.99 55.10 54.21 63.31 +7.44
Mamba-FSCIL (ours) 80.90 76.26 72.97 70.14 67.83 65.74 65.43 64.12 62.31 62.12 61.65 68.13
The comparison with the state-of-the-art methods on miniImageNet, CIFAR-100, and CUB-200 is shown in Table I, Table II, and Table III. The results demonstrate that Mamba-FSCIL consistently outperforms almost all existing static and dynamic adaptation methods by over 1% in terms of last-session accuracy, where all classes are taken into account. Specifically, on miniImageNet, Mamba-FSCIL achieves the highest average accuracy of 69.81% and increases the last-session accuracy to 59.36%, which surpasses the challenging baseline NC-FSCIL [19] by 1.99% and 1.05%, respectively. In CIFAR-100, Mamba-FSCIL maintains the best performance in most sessions, with the accuracy of the last session outperforming NC-FSCIL by 1.4%. Notably, Mamba-FSCIL significantly surpasses dynamic methods such as FeSSSS [89] with an average accuracy improvement of 7.22% over FeSSSS. In the more challenging fine-grained CUB-200 dataset, Mamba-FSCIL leads with an average accuracy of 68.13%, outperforming all static and dynamic methods. Compared to DSN, Mamba-FSCIL shows the remarkable ability to learn complex classification scenarios incrementally, improving the accuracy in the last session by 7.44%, also indicating a lower performance drop over sessions.
To further test Mamba-FSCIL on larger vision models, we evaluated it on the CUB-200 dataset using a Swin Transformer-Tiny [30] backbone pre-trained on ImageNet-1K [117]. We compared it against the state-of-the-art methods using the same backbone, including CLOM [104], NC-FSCIL [19], and Comp-FSCIL [83]. As shown in Table IV, Mamba-FSCIL achieves the highest average session accuracy of 78.55% and the lowest performance drop (PD), reducing PD to 14.00% compared to Comp-FSCIL’s 14.87%. Notably, while Comp-FSCIL leverages human-inspired compositional learning for adaptive concept integration, Mamba-FSCIL, without relying on this extra step, still outperforms Comp-FSCIL, demonstrating the strong potential of our approach.
TABLE IV: FSCIL performance across sessions on CUB-200 using Swin Transformer-Tiny backbone. ”Average Acc.” represents the average accuracy across all sessions. ”PD” indicates the performance drop, calculated as the difference in accuracy between the first and last session. Results for CLOM [104] and Comp-FSCIL [83] are taken from Comp-FSCIL [83], while NC-FSCIL [19] results are from our implementation using the same training settings as Mamba-FSCIL for a fair comparison.
Methods Venue Accuracy in each session (%) $\uparrow$ Average PD
0 1 2 3 4 5 6 7 8 9 10 Acc.
CLOM [104] NeurIPS 2022 86.28 82.85 80.61 77.79 76.34 74.64 73.62 72.82 71.24 71.33 70.50 76.18 15.78
NC-FSCIL [19] ICLR 2023 87.53 84.25 81.72 79.10 77.21 75.52 74.51 74.42 72.26 72.86 72.49 77.44 15.04
Comp-FSCIL [83] ICML 2024 87.67 84.73 83.03 80.04 77.73 75.52 74.32 74.55 73.35 73.15 72.80 77.90 14.87
Mamba-FSCIL (ours) 88.13 85.14 83.41 80.77 77.23 75.73 75.70 75.32 74.18 74.26 74.13 78.55 14.00
@@ 5.3 Ablation Studies
5.3.1 Impact of Individual Component
TABLE V: Ablation study on the impact of dual selective projector and class-sensitive selective scan mechanism on FSCIL performance across different datasets. “Base” indicates the accuracy in the first session; “Final” refers to the last session; “PD” represents the performance drop between the first and last sessions.
Methods Adaptation miniImageNet CIFAR-100 CUB-200
@@ 5.3.1 Impact of Individual Component
We performed ablation studies across the three datasets to assess the impact of our dual selective projector and class-sensitive selective scan mechanism. Utilizing a consistent setup that includes the same backbone, identity branch, memory module, classifier, and classification loss, we explored several configurations: (1) Baseline employs NC-FSCIL with a static MLP-based projection layer, serving as a competitive static adaptation method. (2) + Single Selective SSM Projector integrates a single dynamic SSM projector with one projection branch, updated across all sessions. (3) + Dual Selective SSM Projector uses our dual selective SSM projector with two branches, as introduced in Section 4.2. (4) + Dual Selective SSM Projector + Class-sensitive Selective Scan further integrates our class-sensitive selective scan.
@@
The results in Table V show that the introduction of a single selective SSM projector improves base session accuracy across the three datasets (e.g., from 84.02% to 84.93% on miniImageNet and from 82.52% to 82.80% on CIFAR-100), demonstrating the Mamba-based projector helps to improve the performance in the base session, providing a more powerful feature extractor for the subsequent sessions. However, despite these gains, the final performances on miniImageNet and CIFAR-100 both fall below the baseline, implying that a naive replacement of the MLP projector with the selective SSM projector cannot effectively mitigate catastrophic forgetting, increasing the performance drop as a result. In contrast, when our proposed dual selective SSM projector is adopted, the final performance is also effectively improved over the baseline, reducing the performance drop from 27.04% to 26.01% on miniImageNet and from 26.73% to 26.09% on CIFAR-100. This shows that our dual design is capable of adaptively learning feature shifts of novel classes into their corresponding classifier prototypes with fewer disruptions to the learned knowledge of the base classes. Further improvements are achieved with the integration of our class-sensitive scan mechanism into the dual projector, increasing the last-session accuracy from 56.71% to 57.51% on CIFAR-100 and from 61.06% to 61.65% on CUB-200. These findings validate the critical role of each component in Mamba-FSCIL.
5.3.2 Impact of Losses in Class-sensitive Selective Scan
Figure 3: Effect of suppression and separation losses in Class-sensitive Selective Scan mechanism on feature-classifier alignment and base-novel class separability in the miniImageNet test set: (a) Average cosine similarity between base-class features and their corresponding classifiers; (b) Average cosine similarity between base and novel classes.
We investigated the impacts of the suppression and separation losses introduced in Section 4.3 by analyzing the features extracted from the test set of the miniImageNet. First, we measured the average cosine similarity between the final representations and the classifier prototypes across all base classes in Fig. 3, defined as:
$\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}\cos\langle{\bm{\mu}}_{k,i},\mathbf{w}_{k}\rangle,$ (23)
where ${\bm{\mu}}_{k,i}$ represents the feature vector of sample $i$ in class $k$, $N_{k}$ is the number of samples in class $k$, $\mathbf{w}_{k}$ is the classifier prototype for class $k$, and $K$ is the number of base classes. To assess the separability between base and novel classes, we calculated the average cosine similarity for any possible pairing of base and novel classes feature vectors as:
$\frac{1}{N_{\text{base}}N_{\text{novel}}}\sum_{i=1}^{N_{\text{base}}}\sum_{j=1}^{N_{\text{novel}}}\cos\langle{\bm{\mu}}_{\text{base},i},{\bm{\mu}}_{\text{novel},j}\rangle,$ (24)
where ${\bm{\mu}}_{\text{base},i}$ and ${\bm{\mu}}_{\text{novel},j}$ are feature vectors of the $i$-th base class sample and $j$-th novel class sample, respectively. The terms $N_{\text{base}}$ and $N_{\text{novel}}$ denote the total number of samples in base and novel classes in the test set of miniImageNet.
Fig. 3 (a) shows that the incorporation of suppression loss ${\mathcal{L}}_{\text{supp}}$ helps to improve base-class alignment between features and classifier prototypes in late sessions, facilitating the maintenance of base-class performance. When the separation loss ${\mathcal{L}}_{\text{sep}}$ is further added, as shown in Fig. 3 (b), the similarities of features between the base-class and novel-class are significantly reduced, which leaves sufficient feature space for incremental training sessions and boosts the novel-class accuracy.
5.3.3 Effect of Freezing vs. Training Projection Branches
Figure 4: Impact of freezing vs. training the identity branch $p$ and the selective SSM branch $g^{\text{base}}$ during incremental sessions on miniImageNet: (a) Accuracy for all classes, (b) Accuracy for base classes.
We assessed the impact of freezing the identity branch $p$ and the selective SSM branch $g^{\text{base}}$ from the base session during incremental learning. Models are trained on the miniImageNet dataset, utilizing a ResNet-12 backbone. Two configurations are compared: one with $p$ and $g^{\text{base}}$ frozen (as described in Section 4.2) and the other one with them trainable throughout the incremental sessions. The results in Fig. 4 indicate that the frozen configuration achieves higher accuracy across all classes (Fig. 4 (a)) and specifically for base classes (Fig. 4 (b)) in most sessions. This result implies that freezing $p$ and $g^{\text{base}}$ effectively preserves the base-session knowledge, mitigating catastrophic forgetting.
5.3.4 Sensitivity Analysis of Hyper-parameters $\alpha$ and $\beta$
TABLE VI: Impact of hyper-parameters $\alpha$ and $\beta$ on performance in the miniImageNet dataset. “Average Acc.” represents the average accuracy across all sessions.
@@ 5.3.4 Sensitivity Analysis of Hyper-parameters $\alpha$ and $\beta$
We conducted a sensitivity analysis on the hyper-parameters $\alpha$ and $\beta$ of Eq. (22) in Table VI to evaluate their impact on our model’s average accuracy on miniImageNet. Each experiment varies one parameter with the other fixed. The results show that when $\alpha$ and $\beta$ are set within a proper range, the performance only varies slightly, demonstrating the model’s robustness to the hyper-parameters.
5.3.5 Impact of the Number of Training Samples
Figure 5: Performance across different sessions for 1-shot, 5-shot, 10-shot, and 20-shot learning scenarios on miniImageNet, CIFAR-100, and CUB-200 datasets.
We explored the effect of varying the number of training samples per class, or ”shots,” on Mamba-FSCIL’s performance. As shown in Fig. 5, increasing the number of training examples from 1-shot to 20-shot consistently improves accuracy across all sessions for the miniImageNet, CIFAR-100, and CUB-200 datasets. This demonstrates that Mamba-FSCIL benefits from the additional novel class training samples while effectively preserving base class knowledge, and preventing catastrophic forgetting.
5.3.6 Impact of the Number of Scan Paths
TABLE VII: Impact of the number of scan paths ($K$) on model performance for CIFAR-100 and CUB-200 datasets. “Base” refers to accuracy in the first session; “Final” to the accuracy in the last session; and “Avg” to the average accuracy across all sessions.
K CIFAR-100 CUB-200
[…]
1 82.35 55.94 67.23 80.41 60.84 67.48
2 82.65 57.13 67.53 80.49 61.08 67.86
4 82.80 57.51 68.14 80.90 61.65 68.13
We evaluated the influence of varying the number of scan paths ($K$) on Mamb-FSCIL’s performance on CIFAR-100 and CUB-200 datasets, as described in Table VII. Increasing the scan paths consistently enhances all performance metrics—base, final, and average session accuracies. Specifically, on CIFAR-100, elevating the scan paths from $1$ (top-left to bottom-right) to $2$ (adding bottom-right to top-left) boosts final session accuracy by 1.19%, from 55.94% to 57.13%, increases base session accuracy from 82.35% to 82.65%, and raises average accuracy from 67.23% to 67.53%. Extending to four paths (adding top-right to bottom-left and bottom-left to top-right) further improves final session accuracy to 57.51%, base session accuracy to 82.80%, and average accuracy to 68.14%. Observations on CUB-200 reflect similar trends. These findings highlight the substantial benefits of more scan paths in improving model performance, particularly their role in enhancing the integration of contextual information from multiple directions, which is crucial for handling complex visual tasks in dynamic learning environments. By default, we employed four scan paths in all experiments.
5.4 Comparison with Different Projection Architectures
TABLE VIII: Performance comparison of different projection architectures on CIFAR-100.
[…]
Dual Transformer 82.45 55.05 66.46
Dual MLP 82.57 56.02 67.57
Dual Selective SSM (ours) 82.80 57.24 68.08
To validate the effectiveness of the selective SSM projection branches in the Dual Selective SSM Projector for FSCIL, we performed a comparative experiment on the CIFAR-100 dataset. In this experiment, we replaced the selective SSM branches in our dual-branch architecture with MLP and Transformer structures while maintaining similar parameter counts across all models using a ResNet-12 backbone. The MLP setup includes three layers of multi-layer perceptrons, while the Transformer configuration incorporates a multi-head attention mechanism with 8 heads followed by a feed-forward network. Suppression loss with $\alpha=50$ is applied across all configurations and is directly applied to the output of the newly attached branch ($g^{\text{inc}}$) in the MLP and Transformer-based setups.
As shown in Table VIII, the Dual Selective SSM architecture consistently outperforms both MLP and Transformer setups across all critical metrics: base accuracy, final session accuracy, and average accuracy. Specifically, the Dual Selective SSM architecture enhances final session accuracy by 2.19% compared to the Transformer result and by 1.22% compared to the MLP setup. Furthermore, it achieves an average accuracy improvement of 1.62% over the Transformer and 0.51% over MLP.
5.5 Comparison with Dynamic Adaptation Method
We compared our Mamba-FSCIL with DSN [23] in Fig. 6, particularly in terms of parameter efficiency and session-wise performance. Both methods use the same ResNet18 backbone. As shown in Fig. 6 (a), Unlike DSN, which continually increases its parameter amount to learn new classes, Mamba-FSCIL armed with a dual selective SSM projector dynamically adjusts its parameters without expanding the parameter space. In addition, Fig. 6 (b) illustrates that Mamba-FSCIL consistently outperforms DSN in all sessions, demonstrating Mamba-FSCIL’s superior ability to retain old knowledge while learning new classes.
5.6 Visualization of Activation Maps
Figure 7: Activation maps for base and novel classes on the CUB-200 dataset. (a) Base classes; (b) Novel classes. Activation maps are generated using GradCAM from the $g^{\text{base}}$ and $g^{\text{inc}}$ branches.
We used GradCAM [118] to visualize the activation maps of a Swin Transformer-Tiny model trained on the CUB-200 dataset after 10 incremental sessions. The activation maps from the $g^{\text{base}}$ and $g^{\text{inc}}$ branches are visualized to assess their roles in handling base and novel classes.
Fig. 7 (a) shows that for images from the base classes, the $g^{\text{base}}$ branch significantly influences the final output, with activation regions effectively covering the main subject, indicating that this branch, fixed during incremental training, retains a strong representation of base classes. This confirms its role in preserving learned knowledge. In contrast, the $g^{\text{inc}}$ branch shows minimal activation, reducing interference with established knowledge. For novel classes, as shown in Fig. 7 (b), the fixed $g^{\text{base}}$ branch struggles to represent novel classes adequately, with incomplete activation coverage that could lead to classification errors. However, the $g^{\text{inc}}$ branch compensates by fully covering the subject, enhancing the model’s ability to represent novel classes. These visualizations demonstrate that the $g^{\text{base}}$ branch effectively preserves knowledge from base classes, while the $g^{\text{inc}}$ branch improves representation for novel classes without disrupting existing knowledge.